# On braided simple extensions and braided non-semisimple near-group categories

### v1

@@ Abstract.
Let $\mathcal{C}$ be a finite tensor category with a decomposition $\mathcal{C}\cong\mathcal{D}\oplus{\mathcal{M}}$ into full abelian categories such that $\mathcal{D}$ is a tensor subcategory and $\mathcal{{M}}$ has a single simple projective object. We call $\mathcal{C}$ a simple extension of $\mathcal{D}$. When $\mathcal{C}$ is braided and $\mathcal{D}$ is a braided subcategory, we will call $\mathcal{C}$ a braided simple extension.
Applying the notion of a simple extension to a pointed non-semisimple category $\mathcal{D}$ gives a natural generalization to the family of near-group categories, which was first introduced by [S1]. Braided near-group categories were fully classified in [S2] and [T].
In this paper, we study the structure of braided simple extensions of a braided finite tensor category $\mathcal{D}$ in general, and specifically when $\mathcal{D}$ is pointed. In particular, we classify non-degenerate braided non-semisimple near-group categories, and prove that any braided non-semisimple near-group category $\mathcal{C}$ is ”an extension” of a non-degenerate braided non-semisimple near-group category by $\text{Rep}(G)$, where $G$ is the Picard group of a ”canonical” symmetric subcategory of $\mathcal{C}$.
@@ 1. Introduction
A classical result in the theory of finite tensor categories is that if $\mathcal{D}$ has a fiber functor, then Vec is a semisimple $\mathcal{D}$-bimodule category, and $\mathcal{C}:=\mathcal{D}\oplus\text{Vec}$ is also a $\mathcal{D}$-bimodule category (see [EGNO]). A natural question that arises is: when can we define a monoidal structure on $\mathcal{C}$. To be more precise, we ask:
@@
In fusion categories, there is a family of categories called near-group categories, which were introduced in [S1] (though classification of such categories appeared even earlier in [TY]). A near-group category is a fusion category $\mathcal{C}$ with a unique non-invertible object $Q$. Denoting the group of simple invertible objects of $\mathcal{C}$ by $G=\{L_{i}\}$, then a key parameter in the structure of $\mathcal{C}$ is the parameter $r\in\mathbb{N}$ such that $Q\otimes Q\cong\oplus_{i}L_{i}\oplus rQ$. This near-group category is usually denoted by $\mathcal{C}(G,r)$, and we say that $\mathcal{C}$ has a fusion rule $(G,r)$.
There are many known results regarding near-group categories. For instance, in their work, [TY] proved that every near-group category with fusion rule $(G,0)$ must be non-strict, that is, must have non-trivial associativity constraint (they also determine the structure constants of the possible associativity constraints in such categories). In addition, [S2] classified braided near-group categories with fusion rule $(G,0)$, and their braid structures (including the structure of the underlying group $G$), and [T] classified braided near-group categories with fusion rule $(G,r)$ for $r>0$. These results strongly rely on two factors:
@@
In this paper, we wish to take a first step toward classifing non-semi-simple near-group categories, by focusing on braided non-semisimple near-group categories, i.e. braided simple extensions of braided pointed finite tensor categories, admitting a fiber functor. Such a classification should be easier, due to the classification of braided pointed finite tensor categories with a fiber functor, which is given in [BN1].
Formally, in this work we consider triplets of simple extensions $(\mathcal{C},\mathcal{D},\mathcal{M})$ for which $\mathcal{D}$ is braided, $\mathcal{M}$ is equivalent as a $\mathcal{D}$-bimodule category to Vec and $\mathcal{C}$ is the simple extension of $\mathcal{D}$, such that $\mathcal{C}$ is braided and $\mathcal{D}$ is a braided tensor subcategory of $\mathcal{C}$. An example for such a category is given in [GS]. On the other hand, there are braided finite tensor categories with simple extensions that are not braided - for instance $\text{Rep}(H_{4})$ is braided but $\text{Rep}(A^{\prime\prime}_{C_{4}})$ is not, because $A^{\prime\prime}_{C_{4}}$ has no quasi-triangular structure (see [M]).
In the fusion case, [T] proves that there are braided near-group categories with fusion rule $(G,r)$ for $r>0$. In this paper, we prove that this fails in non-semisimple near-group categories. That is, we prove the following theorem:
@@
Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$. In particular, a non-semisimple braided near-group category is weakly integral.
In their work ([BN2], [BN3]), the authors defined the notion of a central exact sequence of finite tensor categories as an exact sequence $\mathcal{E}\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}$ of finite tensor categories with exact tensor functors $i:\mathcal{E}\to\mathcal{C}$ and $F:\mathcal{C}\to\mathcal{D}$ such that $i$ is a full embedding, $F$ is surjective and normal, meaning that for every $Y\in\mathcal{C}$ there exists a subobject $X^{\prime}\in\mathcal{C}$ such that $F(X^{\prime})$ is the maximal subobject of $F(X)$ that is a direct sum of copies of $\mathbbm{1}_{\mathcal{D}}$, and $F\circ i$ is a fiber functor of $\mathcal{E}$. Such an exact sequence is called braided if the categories $\mathcal{E},\mathcal{C},\mathcal{D}$ and the functors $i,F$ are braided, and it is called a modularization exact sequence if $\mathcal{D}$ is modular (i.e, non-degenerate). In this paper, we prove that braided non-semisimple near-group categories are ”extensions” of non-degenerate braided non-semisimple near-group categories by symmetric fusion categories. Formally, we prove:
@@
In their work, [GS] classified non-degenerate braided finite tensor categories with a Lagrangian subcategory braid equivalent to $\textrm{sRep}(W)$ (the category of supermodules of the exterior algebra over a purely odd supervector space $W$), and they proved that some of these categories are simple braided extensions of $\textrm{sRep}(W\oplus W^{*})$ (though they do not use the term ”braided simple extension”). In this paper, we prove that these examples are, in a way, unique. To be exact, we prove:
@@
This result is a slightly weaker version of the result given in [S2] concerning braided (fusion) near-group categories. In particular, this Theorem gives a categorical proof to the result of [M] stated above that there is no braided simple extension to $\text{Rep}(H_{4})$, in $\text{Rep}(H_{4})$ we have $n_{g}+n_{p}=2$.
@@ 2.2. Braiding, symmetry, twists and Müger centralizer
for $a$ the associativity constraint of $\mathcal{C}$. For more details, see for instance [EGNO].
For a braided finite tensor category $\mathcal{C}$, let $s$ be the squared braiding, i.e., $s_{X,Y}:=c_{Y,X}\circ c_{X,Y}$ for every $X,Y\in\mathcal{C}$. Recall that two objects $X,Y\in{\mathcal{C}}$ centralize each other if $s_{X,Y}=\text{id}_{X\otimes{Y}}$, and that the centralizer $\mathcal{D}^{\prime}$ of a full tensor subcategory $\mathcal{D}\subseteq\mathcal{C}$ is the full subcategory of $\mathcal{C}$ consisting of all objects which centralize every object of $\mathcal{D}$ (see, e.g., [DGNO]). The notion of Müger’s centralizer enables us to distinguish between different types of braided finite tensor categories. For finite braided tensor categories $\mathcal{D}\subset\mathcal{C}$ we say that:
@@
By [Sh], we have in general for $\mathcal{D}\subset\mathcal{C}$
@@ 2.3. De-equivariantization
In this subsection, we briefly recall the notion of de-equivariantization and some results which we will use in the classification of braided non-semisimple near-group categories. For more details, see [EGNO].
@@
for $\bar{Y}$ the image of $Y$ under the forgetful functor $\textrm{Mod}_{\mathcal{C}}(A)\to\mathcal{C}$ (see Lemma 7.8.12 in [EGNO]).
It is well known that for a finite group $G$, $A_{G}:=\text{Fun}(G,\mathbbm{k})$ is an algebra object in the category $\text{Rep}(G)$, see for instance [EGNO].
@@
The proof is parallel to Proposition 8.23.8 in [EGNO], using the inclusion as a fully faithful braided tensor functor $\text{Rep}(G)\to\mathcal{C}^{\prime}$. ∎
@@
In the case of fusion categories, if $\mathcal{C}$ has a $G$-de-equivariantization, then we have $(\mathcal{C}_{G})^{G}\simeq\mathcal{C}$, for $\mathcal{D}^{G}$ the $G$-equivariantization of $\mathcal{D}$ by $G$ (see [EGNO] for details on $G$-equivariantization). In finite tensor categories, this is not always true. For example, for $H$ the Sweedler’s Hopf algebra, $\text{Rep}(\mathbb{Z}_{2})$ is central in $\text{Rep}(H)$, and $\text{Rep}(H)_{\mathbb{Z}_{2}}=\text{Vec}$, so $(\text{Rep}(H)_{\mathbb{Z}_{2}})^{\mathbb{Z}_{2}}\not\simeq\text{Rep}(H)$. This happens because the fiber functor $F:\text{Rep}(H)\to\text{Vec}$ is not a braided fiber functor.
@@
The free module functor is a surjective tensor functor, so by [EO] it preserves Frobenious-Perron dimensions, thus $\text{\rm FPdim}_{\mathcal{C}_{G}}(X\otimes A)=1$, hence by [EGNO], $X\otimes A$ is invertible in $\mathcal{C}_{G}$. ∎
@@ 2.4. Braided exact sequences of tensor categories
In this subsection, we recall the notion of braided exact sequences of finite tensor categories, which was introduced in [BN3].
@@
Notice that this definition requires in particular that $\mathcal{C}_{1}$ admits a fiber functor. The notion of exact sequences was generalized to avoid this restriction by [EG], but in this work, we will consider exact sequences as defined in [BN3].
@@ 3.1. Basic properties of simple extensions
The category $\mathcal{D}$ has a simple extension, hence by definition has a fiber functor $\mathcal{D}\to\text{Vec}$. In particular, $\mathcal{D}$ is integral, and $\text{\rm FPdim}_{\mathcal{D}}(L)=\dim_{\mathbbm{k}}(L)$ (see [EGNO]). $\mathcal{M}=<Q>$ is a $\mathcal{D}$-module category, and $Q$ is projective, hence we have an isomorphism $L\otimes{Q}\cong{\oplus_{i=1}^{\dim_{\mathbbm{k}}(L)}Q}$, which gives
@@ 3.2. Degeneracy in braided simple extensions
Assume we have $X,Y,Z\in\mathcal{C}$ such that $s_{X\otimes Y}=\text{id}_{X\otimes Y}$ and $\psi:X\simeq Z$, then $s_{Z\otimes Y}=\text{id}_{Z\otimes Y}$, by the following diagram:
$X\otimes Y$$X\otimes Y$$Z\otimes Y$$Z\otimes Y$$\scriptstyle{s_{X\otimes Y}}$$\scriptstyle{\psi\otimes\text{id}_{Y}}$$\scriptstyle{\psi\otimes\text{id}_{Y}}$$\scriptstyle{s_{Z\otimes Y}}$
which commutes due to the naturality of the braiding $c$.
@@
By Corollary 8.10.8 in [EGNO], the distinguished invertible object of $\mathcal{C}$ lies in $\mathcal{E}_{+}$.
@@
The category $\mathcal{C}$ is a symmetric finite tensor category over an algebraically closed field of characteristic 0, thus by [D], $\mathcal{C}$ is equivalent to a category of representations of some finite supergroup, but such a category has no simple projective object, unless it is semisimple itself, thus contradicting our assumption. ∎
@@ 4. Braided simple extensions of pointed tensor categories
In this section, we assume that $\mathcal{D}$ is a pointed braided finite tensor category, and $\mathcal{C}$ is a braided simple extension of $\mathcal{D}$. Concrete examples of such categories can be found in Appendix A in [GS]. Denote the simple projective object in $\mathcal{C}$ by $Q$. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a pointed braided fusion subcategory, which we denote by $\mathcal{D}_{f}$. This category corresponds to a pair $(G,q)$ for $G$ a finite abelian group and $q$ a quadratic form on $G$. The group $G$ is simply $\text{Pic}(\mathcal{D})$ (the Picard group), and the squared braiding $s_{L_{g},L_{h}}:L_{g}\otimes L_{h}\to L_{h}\otimes L_{g}$ of two simple objects $L_{g},L_{h}\in\mathcal{D}_{f}$ is multiplication by $b(g,h)=\frac{q(gh)}{q(g)q(h)}$.

### v2

@@ Abstract.
We study simple extensions of pointed finite tensor categories, that is, tensor categories $\mathcal{C}$ admitting an abelian decomposition $\mathcal{C}\cong\mathcal{D}\oplus\mathcal{M}$ where $\mathcal{D}$ is a pointed tensor subcategory and $\mathcal{M}$ has a unique simple projective object. Such categories provide a natural generalization of near-group categories. Our results concern the braided case.
We prove that every non-degenerate braided non-semisimple near-group category is a braided simple extension of $\textrm{sRep}(W\oplus W^{*})$ with non-trivial braiding for which $\textrm{sRep}(W)$ is Lagrangian. Moreover, any braided non-semisimple near-group category $\mathcal{C}$ arises canonically as an extension of such a category by $\text{Rep}(G)$, where $G$ is the Picard group of a symmetric subcategory determined by the unique simple projective object of $\mathcal{C}$.
Keywords: finite tensor categories, near-group categories, non-semisimple categories, braided tensor categories
MSC 2020: 18M05, 18M15.
@@ 1. Introduction
A classical result in the theory of finite tensor categories is that if $\mathcal{D}$ has a fiber functor, then Vec is a semisimple $\mathcal{D}$-bimodule category, and $\mathcal{C}:=\mathcal{D}\oplus\text{Vec}$ is also a $\mathcal{D}$-bimodule category (see [1]). A natural question that arises is: when can we define a monoidal structure on $\mathcal{C}$. To be more precise, we ask:
@@
In fusion categories, there is a family of categories called near-group categories, which were introduced in [2] (though classification of such categories appeared even earlier in [3]). A near-group category is a fusion category $\mathcal{C}$ with a unique non-invertible object $Q$. Denoting the group of simple invertible objects of $\mathcal{C}$ by $G=\{L_{i}\}$, then a key parameter in the structure of $\mathcal{C}$ is the parameter $r\in\mathbb{N}$ such that $Q\otimes Q\cong\oplus_{i}L_{i}\oplus rQ$. This near-group category is usually denoted by $\mathcal{C}(G,r)$, and we say that $\mathcal{C}$ has a fusion rule $(G,r)$.
There are many known results regarding near-group categories. For instance, in their work, [3] proved that every near-group category with fusion rule $(G,0)$ must be non-strict, that is, must have non-trivial associativity constraint (they also determine the structure constants of the possible associativity constraints in such categories). In addition, [4] classified braided near-group categories with fusion rule $(G,0)$, and their braid structures (including the structure of the underlying group $G$), and [5] classified braided near-group categories with fusion rule $(G,r)$ for $r>0$. These results strongly rely on two factors:
@@
In this paper, we wish to take a first step toward classifying non-semi-simple near-group categories, by focusing on braided non-semisimple near-group categories, i.e. braided simple extensions of braided pointed finite tensor categories, admitting a fiber functor. Such a classification should be easier, due to the classification of braided pointed finite tensor categories with a fiber functor, which is given in [7].
Formally, in this work we consider triplets of simple extensions $(\mathcal{C},\mathcal{D},\mathcal{M})$ for which $\mathcal{D}$ is braided, $\mathcal{M}$ is equivalent as a $\mathcal{D}$-bimodule category to Vec and $\mathcal{C}$ is the simple extension of $\mathcal{D}$, such that $\mathcal{C}$ is braided and $\mathcal{D}$ is a braided tensor subcategory of $\mathcal{C}$. An example for such a category is given in [8]. On the other hand, there are braided finite tensor categories with simple extensions that are not braided - for instance $\text{Rep}(H_{4})$ is braided but $\text{Rep}(A^{\prime\prime}_{C_{4}})$ is not, because $A^{\prime\prime}_{C_{4}}$ has no quasi-triangular structure (see [6]).
In the fusion case, [5] proves that there are braided near-group categories with fusion rule $(G,r)$ for $r>0$. In this paper, we prove that this fails in non-semisimple near-group categories. That is, we prove the following theorem:
@@
Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$, and consequently $\mathcal{C}$ is weakly integral.
In their work ([9], [10]), the authors defined the notion of a central exact sequence of finite tensor categories as an exact sequence $\mathcal{E}\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}$ of finite tensor categories with exact tensor functors $i:\mathcal{E}\to\mathcal{C}$ and $F:\mathcal{C}\to\mathcal{D}$ such that $i$ is a full embedding, $F$ is surjective and normal, meaning that for every $Y\in\mathcal{C}$ there exists a subobject $X^{\prime}\in\mathcal{C}$ such that $F(X^{\prime})$ is the maximal subobject of $F(X)$ that is a direct sum of copies of $\mathbbm{1}_{\mathcal{D}}$, and $F\circ i$ is a fiber functor of $\mathcal{E}$. Such an exact sequence is called braided if the categories $\mathcal{E},\mathcal{C},\mathcal{D}$ and the functors $i,F$ are braided, and it is called a modularization exact sequence if $\mathcal{D}$ is modular (i.e, non-degenerate). In this paper, we prove that braided non-semisimple near-group categories are ”extensions” of non-degenerate braided non-semisimple near-group categories by symmetric fusion categories. Formally, we prove:
@@
In their work, [8] classified non-degenerate braided finite tensor categories with a Lagrangian subcategory braid equivalent to $\textrm{sRep}(W)$ (the category of supermodules of the exterior algebra over a purely odd supervector space $W$), and they proved that some of these categories are simple braided extensions of $\textrm{sRep}(W\oplus W^{*})$ (though they do not use the term ”braided simple extension”). In this paper, we prove that these examples are, in a way, unique. To be exact, we prove:
@@
This result is a slightly weaker version of the result given in [4] concerning braided (fusion) near-group categories. In particular, this Theorem gives a categorical proof to the result of [6] stated above that there is no braided simple extension to $\text{Rep}(H_{4})$, in $\text{Rep}(H_{4})$ we have $n_{g}+n_{p}=2$.
@@ 2.2. Braiding, symmetry, twists and Müger centralizer
for $a$ the associativity constraint of $\mathcal{C}$. For more details, see for instance [1].
For a braided finite tensor category $\mathcal{C}$, let $s$ be the squared braiding, i.e., $s_{X,Y}:=c_{Y,X}\circ c_{X,Y}$ for every $X,Y\in\mathcal{C}$. Recall that two objects $X,Y\in{\mathcal{C}}$ centralize each other if $s_{X,Y}=\text{id}_{X\otimes{Y}}$, and that the centralizer $\mathcal{D}^{\prime}$ of a full tensor subcategory $\mathcal{D}\subseteq\mathcal{C}$ is the full subcategory of $\mathcal{C}$ consisting of all objects which centralize every object of $\mathcal{D}$ (see, e.g., [11]). The notion of Müger’s centralizer enables us to distinguish between different types of braided finite tensor categories. For finite braided tensor categories $\mathcal{D}\subset\mathcal{C}$ we say that:
@@
By [12], we have in general for $\mathcal{D}\subset\mathcal{C}$
@@ 2.3. De-equivariantization
In this subsection, we briefly recall the notion of de-equivariantization and some results which we will use in the classification of braided non-semisimple near-group categories. For more details, see [1].
@@
for $\bar{Y}$ the image of $Y$ under the forgetful functor $\textrm{Mod}_{\mathcal{C}}(A)\to\mathcal{C}$ (see Lemma 7.8.12 in [1]).
It is well known that for a finite group $G$, $A_{G}:=\text{Fun}(G,\mathbbm{k})$ is an algebra object in the category $\text{Rep}(G)$, see for instance [1].
@@
The proof is parallel to Proposition 8.23.8 in [1], using the inclusion as a fully faithful braided tensor functor $\text{Rep}(G)\to\mathcal{C}^{\prime}$. ∎
@@
In the case of fusion categories, if $\mathcal{C}$ has a $G$-de-equivariantization, then we have $(\mathcal{C}_{G})^{G}\simeq\mathcal{C}$, for $\mathcal{D}^{G}$ the $G$-equivariantization of $\mathcal{D}$ by $G$ (see [1] for details on $G$-equivariantization). In finite tensor categories, this is not always true. For example, for $H$ the Sweedler’s Hopf algebra, $\text{Rep}(\mathbb{Z}_{2})$ is central in $\text{Rep}(H)$, and $\text{Rep}(H)_{\mathbb{Z}_{2}}=\text{Vec}$, so $(\text{Rep}(H)_{\mathbb{Z}_{2}})^{\mathbb{Z}_{2}}\not\simeq\text{Rep}(H)$. This happens because the fiber functor $F:\text{Rep}(H)\to\text{Vec}$ is not a braided fiber functor.
@@
The free module functor is a surjective tensor functor, so by [13] it preserves Frobenius-Perron dimensions, thus $\text{\rm FPdim}_{\mathcal{C}_{G}}(X\otimes A)=1$, hence by [1], $X\otimes A$ is invertible in $\mathcal{C}_{G}$. ∎
@@ 2.4. Braided exact sequences of tensor categories
In this subsection, we recall the notion of braided exact sequences of finite tensor categories, which was introduced in [10].
@@
Notice that this definition requires in particular that $\mathcal{C}_{1}$ admits a fiber functor. The notion of exact sequences was generalized to avoid this restriction by [14], but in this work, we will consider exact sequences as defined in [10].
@@ 3.1. Basic properties of simple extensions
The category $\mathcal{D}$ has a simple extension, hence by definition has a fiber functor $\mathcal{D}\to\text{Vec}$. In particular, $\mathcal{D}$ is integral, and $\text{\rm FPdim}_{\mathcal{D}}(L)=\dim_{\mathbbm{k}}(L)$ (see [1]). $\mathcal{M}=<Q>$ is a $\mathcal{D}$-module category, and $Q$ is projective, hence we have an isomorphism $L\otimes{Q}\cong{\oplus_{i=1}^{\dim_{\mathbbm{k}}(L)}Q}$, which gives
@@ 3.2. Degeneracy in braided simple extensions
Assume we have $X,Y,Z\in\mathcal{C}$ such that $c_{X,Y}=\lambda\text{id}_{Y\otimes X}$ for some $\lambda\in\mathbbm{k}^{*}$ and $\psi:X\simeq Z$, then $c_{Z,Y}=\lambda\text{id}_{Y\otimes Z}$, by the following diagram:
$X\otimes Y$$Y\otimes X$$Z\otimes Y$$Y\otimes Z$$\scriptstyle{c_{X,Y}}$$\scriptstyle{\psi\otimes\text{id}_{Y}}$$\scriptstyle{\text{id}_{Y}\otimes\psi}$$\scriptstyle{c_{Z,Y}}$
which commutes due to the naturality of the braiding $c$. The same holds for $s_{X,Y}$ and $s_{Z,Y}$.
@@
By Corollary 8.10.8 in [1], the distinguished invertible object of $\mathcal{C}$ lies in $\mathcal{E}_{+}$.
@@
The category $\mathcal{C}$ is a symmetric finite tensor category over an algebraically closed field of characteristic 0, thus by [15], $\mathcal{C}$ is equivalent to a category of representations of some finite supergroup, but such a category has no simple projective object, unless it is semisimple itself, thus contradicting our assumption. ∎
@@ 4. Braided simple extensions of pointed tensor categories
In this section, we assume that $\mathcal{D}$ is a pointed braided finite tensor category, and $\mathcal{C}$ is a braided simple extension of $\mathcal{D}$. Concrete examples of such categories can be found in Appendix A in [8]. Denote the simple projective object in $\mathcal{C}$ by $Q$. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a pointed braided fusion subcategory, which we denote by $\mathcal{D}_{f}$. This category corresponds to a pair $(G,q)$ for $G$ a finite abelian group and $q$ a quadratic form on $G$. The group $G$ is simply $\text{Pic}(\mathcal{D})$ (the Picard group), and the squared braiding $s_{L_{g},L_{h}}:L_{g}\otimes L_{h}\to L_{h}\otimes L_{g}$ of two simple objects $L_{g},L_{h}\in\mathcal{D}_{f}$ is multiplication by $b(g,h)=\frac{q(gh)}{q(g)q(h)}$.